OpenStudy (anonymous):

Jarod is having a problem with rabbits getting into his vegetable garden, so he decides to fence it in. The length of the garden is 7 feet more than 6 times the width. He needs 70 feet of fencing to do the job. Find length of the garden.

OpenStudy (anonymous):

First create a variable for the width... lets call it 'w'. Then... Length = 6w + 7 Width = w We can find w because we know that the total fencing is 70 ft. So... 70ft = 2*(6w+7) + 2*(w) If you solve this you will find that w is equal to 4 feet. You can then plug in w to the equation for the length of the garden... Length = 6(4) + 7 Length = 31 Feet